OpenStudy (anonymous):

7/4x+7/12=1/3x-1/6. Solve. PLease explain how to find answer.

OpenStudy (anonymous):

\[\frac{ 7 }{ 4 }x+\frac{ 7 }{ 12 }=\frac{ 1 }{ 3 }x-6\]

OpenStudy (anonymous):

Sorry, thats \[\frac{ 7 }{ 4 }x+\frac{ 7 }{ 12 }=\frac{ 1 }{ 3 }x-\frac{ 1 }{ 6 }\]

OpenStudy (anonymous):

Anyone?